# LLMatic: Neural Architecture Search via Large Language Models and Quality Diversity Optimization

### v7

@@ 5. Experiments on NAS-bench-201
Next, we extend our experimentation of LLMatic on NAS-bench-201 benchmark (Dong and Yang, 2020), which searches a cell block for a constant neural network structure. The structure is initiated with one 3-by-3 convolution with 16 output channels and a batch normalization layer (Ioffe and Szegedy, 2015). The main body of the skeleton includes three stacks of cells, connected by a residual block. Each cell is stacked 5 times, with the number of output channels as 16, 32 and 64 for the first, second and third stages, respectively. The intermediate residual block is the basic residual block with a stride of 2 (He et al., 2016), which serves to downsample the spatial size and double the channels of an input feature map. The shortcut path in this residual block consists of a 2-by-2 average pooling layer with a stride of 2 and a 1-by-1 convolution. The skeleton ends with a global average pooling layer to flatten the feature map into a feature vector. Classification uses a fully connected layer with a softmax layer to transform the feature vector into the final prediction.
The specified cell within the search domain is depicted as a densely connected directed acyclic graph with four nodes and six edges; here, nodes symbolise feature maps while edges denote operations. There are five possible operations: (1) zeroize, (2) skip connection, (3) 1-by-1 convolution, (4) 3-by-3 convolution, and (5) 3-by-3 average pooling layer. Zeroize drops out the associated edge operation. Given five operations to choose from, the aggregate count of potential search spaces comes to $156255^{6}=$ 15625 cell combinations. Evaluations are carried out on CIFAR10, CIFAR100 (Krizhevsky et al., 2009), and ImageNet16-120 (Chrabaszcz et al., 2017). ImageNet16-120 is a variant of ImageNet dataset (Russakovsky et al., 2015) which is downsampled to 16x16 image sizes and contains the first 120 classes.
@@ 5.1. Results
Table 1. A comparison of test accuracy on NAS-bench-201 LLMatic with $\Lambda$-DARTS (Movahedi et al., 2022), having the most near-to-optimal result, GENIUS (Zheng et al., 2023), a GPT-4 based neural architecture search algorithm, Random Search implemented by NAS-bench-201 authors (Dong and Yang, 2020) and DARTS (Liu et al., 2018). We provide optimal accuracy for reference, which is the maximum accuracy that can be achieved in NAS-bench-201. LLMatic achieves near-to-optimal results but has more variation than the comparing methods.
@@
We compare our results with GPT-4 based NAS algorithm GENIUS (Zheng et al., 2023) as an LLM baseline, $\Lambda$-DARTS (Movahedi et al., 2022) as it achieves close to optimal result, where optimal is the maximum test accuracy, and Random Search. As Table 1 indicates, LLMatic has better results than simple GPT-4 based NAS and close to the state-of-the-art and optimal results.
Furthermore, in Figure 5 we look into the found networks by LLMatic over each generation. We observe that the distribution of found networks is spread wide in the search space. This is due to the procedural nature and exploration capabilities of LLMatic through the prompt archive. To demonstrate near-to-optimal networks we look into Table 2 for maximum ranked networks based on test accuracies searched by LLMatic.
@@ 6. Conclusion and Future Work
To conclude, we present LLMatic: a novel Neural Architecture Search (NAS) algorithm that harnesses the power of Large Language Models (LLMs) and Quality-Diversity (QD) optimization algorithms. LLMatic successfully finds competitive networks that are diverse in architecture. We show empirically that LLMatic can find more than 20 competitive networks in CIFAR-10 and near-to-optimal networks in NAS-bench-201, using only 2000 evaluations. LLMatic decreases the max population size per generation to only 100. LLMatic achieves this while relying on a 6.1B parameter language model. Furthermore, we show that each component in LLMatic is necessary. We do an extensive ablation study and find that LLMatic finds the network with the best accuracy among other variants.
LLMatic achieves this with many constraints in hand. Firstly, we use CodeGen-6.1B code generation LLM, which is a smaller language model when compared to existing LLMs. This demonstrates how computationally efficient LLMatic is, and gives us reason to believe that further gains could be unlocked by incorporating larger language models Secondly, due to computational resources, we keep our searches to $2000$, and still find competitive networks.

### v8

@@ 5. Experiments on NAS-bench-201
Next, we extend our experimentation of LLMatic to the NAS-bench-201 benchmark (Dong and Yang, 2020), which searches a cell block for a constant neural network structure. The structure is initiated with one $3\times 3$ convolution with $16$ output channels and a batch normalization layer (Ioffe and Szegedy, 2015). The main body of the skeleton includes three stacks of cells, connected by a residual block. Each cell is stacked 5 times, with the number of output channels as $16$, $32$ and $64$ for the first, second and third stages, respectively. The intermediate residual block is the basic residual block with a stride of $2$ (He et al., 2016), which serves to downsample the spatial size and double the channels of an input feature map. The shortcut path in this residual block consists of a $2\times 2$ average pooling layer with a stride of $2$ and a $1\times 1$ convolution. The skeleton ends with a global average pooling layer to flatten the feature map into a feature vector. Classification uses a fully connected layer with a softmax layer to transform the feature vector into the final prediction.
The specified cell within the search domain is depicted as a densely connected directed acyclic graph with four nodes and six edges; here, nodes symbolise feature maps while edges denote operations. There are five possible operations: (1) zeroize, (2) skip connection, (3) $1\times 1$ convolution, (4) $3\times 3$ convolution, and (5) $3\times 3$ average pooling layer. Zeroize drops out the associated edge operation. Given five operations to choose from, the aggregate count of potential search spaces is $156255^{6}=$ 15625 cell combinations. Evaluations are carried out on CIFAR10, CIFAR100 (Krizhevsky et al., 2009), and ImageNet16-120 (Chrabaszcz et al., 2017). ImageNet16-120 is a variant of ImageNet dataset (Russakovsky et al., 2015) which is downsampled to 16x16 image sizes and contains the first 120 classes.
@@ 5.1. Results
Table 1. A comparison of test accuracy on the NAS-bench-201 benchmark. We provide the optimal accuracy for reference, which is the maximum accuracy that can be achieved in NAS-bench-201. The results for LLMatic are averaged over 10 runs.
@@
We compare our results with the GPT-4-based NAS algorithm GENIUS (Zheng et al., 2023), which serves as an LLM baseline, as well as random search. We also compare to prior work, including DARTS  (Liu et al., 2018) and $\Lambda$-DARTS (Movahedi et al., 2022), which achieves near-optimal results. As Table 1 indicates, LLMatic outperforms the GPT-4-based NAS, and produces results that are near state-of-the-art, although it has more variation than the competing methods.
Furthermore, in Figure 5 we investigate the networks discovered by LLMatic over each generation. We observe that the distribution of found networks is spread wide in the search space. This is due to the procedural nature and exploration capabilities of LLMatic through the prompt archive. To demonstrate near-to-optimal networks we illustrate in Table 2 the maximum ranked networks based on test accuracies searched by LLMatic.
@@ 6. Conclusion and Future Work
To conclude, we present LLMatic: a novel neural architecture search (NAS) algorithm that harnesses the power of large language models (LLMs) and Quality-Diversity (QD) optimization algorithms. LLMatic successfully finds competitive networks that are diverse in architecture. We show empirically that LLMatic can find more than 20 competitive networks in CIFAR-10 and near-to-optimal networks in NAS-bench-201, using only 2000 evaluations. LLMatic decreases the max population size per generation to only 100. LLMatic achieves this while relying on a 6.1B parameter language model. Furthermore, we show that each component in LLMatic is necessary. We conducted an extensive ablation study and found that LLMatic finds the network with the best accuracy among other variants.
LLMatic achieves this with many constraints in hand. Firstly, we use CodeGen-6.1B code generation LLM, which is a smaller language model when compared to existing LLMs. This demonstrates the computationally efficiency of LLMatic, and gives us reason to believe that further gains could be unlocked by incorporating larger language models Secondly, due to computational resources, we keep our searches to $2000$, and still find competitive networks.
@@
7. Acknowledgments
The authors acknowledge the Centre for High Performance Computing (CHPC), South Africa, for providing computational resources to this research project.